OpenStudy (anonymous):

integrate (sqrt((x-1)/x^5)) using substitution (no trig substitution)

OpenStudy (blockcolder):

Trig substitution is not allowed, but I suppose I can do this: \[\int \frac{\sqrt{x-1}}{x^5} \text{d}x\\ \begin{align} \text{Let }& u =\sqrt{x-1}\\ &u^2=x-1\\ &2u du=dx\\ &(u^2+1)^5=x^5 \end{align}\] Thus the integral becomes: \[\int \frac{2u^3}{(u^2+1)^5}\text{d}u\] Now let v=u^2+1.

OpenStudy (anonymous):

wait wait i think u got the question wrongly. Its \[\int\limits\sqrt{x-1\div x^{5}}\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{x^2} \sqrt{\frac{x-1}{x} } dx\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{x^2} \sqrt{1-\frac{1}{x}} dx\]

myininaya (myininaya):

let u be the inside of that square root thingy

OpenStudy (anonymous):

okay i got it until there, and then?

myininaya (myininaya):

\[u=1-\frac{1}{x} => du=\frac{1}{x^2} dx\]

myininaya (myininaya):

\[\int\limits_{}^{}\sqrt{u} du\]

myininaya (myininaya):

You can integrate that expression?

OpenStudy (anonymous):

yessss. thank you! that helps alot! :)